# Optimal upper bounds for anti-canonical volumes of singular toric Fano varieties

### v1

@@ 1.1. Upper bounds for anti-canonical volumes of Fano varieties
and the equality holds if $X\simeq\mathbb{P}(\frac{2u_{d,q}}{1+u_{1,q}},\cdots,\frac{2u_{d,q}}{1+u_{d-1,q}},1,1)$.
@@
1.2. Upper bounds for dual volumes of Fano polytopes
@@
Fix two integers $d\geq 3,q\geq 2$. Let $P$ be a $d$-dimensional $\frac{1}{q}$-lc Fano polytope. Then
@@
and the equality holds if $P^{*}\simeq\operatorname{conv}(\frac{1+u_{1,q}}{q}e_{1}^{*},\cdots,\frac{1+u_{d-1,q}}{q}e_{d-1}^{*},\pm\frac{u_{d,q}}{q}e_{d}^{*}),$ where $e_{i}^{*}$ are the dual basis of the standard basis of $\mathbb{R}^{d}$.
@@ 1.3. Upper bounds for volumes of lattice simplices
At the same time, we can get the following optimal upper bound for the volumes of lattice simplices $S\subset{\mathbf{N}}_{\mathbb{R}}$ such that $\operatorname{int}(\frac{1}{q}S)\cap{\mathbf{N}}=\{{\mathbf{0}}\}$, where ${\mathbf{N}}$ is a lattice of rank $d$ (See Subsection 2.1 below).
@@
Fix two integers $d\geq 3,q>0$. Let $S$ be a $d$-dimensional latttice simplex in ${\mathbf{N}}_{\mathbb{R}}$ such that $\operatorname{int}(\frac{1}{q}S)\cap{\mathbf{N}}=\{{\mathbf{0}}\}$. Then
$\operatorname{vol}(S)\leq\frac{2u_{d,q}^{2}}{d!q},$
@@ 1.4. Organization of this paper
In Section 2, we collect the necessary material from geometry of numbers and toric geometry, introducing the relevant polytopes, the way to bound the volume of a simplex (see Theorem 2.3) and the decomposition strategy to deal with the dual volume of certain polytopes (see Theorem 2.7); In Section 3, we introduce generalized Product-Sum inequalities (see Thereom 3.1) for barycentric coordinates of a $\frac{1}{q}$-lc Fano simplex and use it to characterize when the products of barycentric coordinates attain the minimal ones (see Theorems 3.8 and 3.9); In Section 4, based on the results in Section 3, we apply Corollary 2.4 to get an upper bound for the dual normalized volume of any $d$-dimensional $\frac{1}{q}$-lc Fano simplex when $d\geq 2$, and hence an upper bound for the anti-canonical volumes of $d$-dimensional $\frac{1}{q}$-lc toric Fano varieties of Picard number one (see Theorem 4.2). We construct examples to show that these upper bounds are optimal when $d\geq 3$ (see Example 4.3); In Section 5, we use the decomposition strategy and integration method from [5] to find upper bounds for the dual normalized volumes of $\frac{1}{q}$-lc Fano polytopes; In Section 6, we prove Theorem 1.1, Theorem 1.2 and Theorem 1.3.
@@ 2.1. Geometry of numbers
We refer to [10] for the basic knowledge in geometry of numbers. Fix a positive integer $d$. Let ${\mathbf{N}}\simeq{\mathbb{Z}}^{d}$ be a lattice of rank $d$ and ${\mathbf{M}}$ be the dual lattice of ${\mathbf{N}}$. Denote ${\mathbf{N}}_{\mathbb{R}}:={\mathbf{N}}\otimes_{{\mathbb{Z}}}\mathbb{R}$ and ${\mathbf{M}}_{\mathbb{R}}:={\mathbf{M}}\otimes_{{\mathbb{Z}}}\mathbb{R}$. A polytope $P$ is a convex bounded subset of ${\mathbf{N}}_{\mathbb{R}}$ such that $P$ is the intersection of finitely many affine half-spaces. For a polytope $P$, we use $\operatorname{int}(P)$ and $\operatorname{vert}(P)$ to denote the interior of $P$ and the set of vertices of $P$ respectively. The dual of a polytope $P$ is defined as
$P^{*}:=\{m\in{\mathbf{M}}_{\mathbb{R}}|<m,v>+1\geq 0,\forall v\in P\}.$
@@
A polytope $P\subset{\mathbf{N}}_{\mathbb{R}}$ is called a lattice polytope if $\operatorname{vert}(P)\subset{\mathbf{N}}$. We say $P\subset{\mathbf{N}}_{\mathbb{R}}$ is a Fano polytope if it is a lattice polytope with ${\mathbf{0}}\in\operatorname{int}(P)$ and each vetex $v\in\operatorname{vert}(P)$ is a primitive lattice point. A Fano polytope $P$ is called an $\epsilon$-lc Fano polytope for some $\epsilon>0$ if $\operatorname{int}(\epsilon P)\cap{\mathbf{N}}=\{{\mathbf{0}}\}$. Here ${\mathbf{0}}$ denotes the origin of ${\mathbf{N}}_{\mathbb{R}}$.
@@
A simplex $S\subset{\mathbf{N}}_{\mathbb{R}}$ is a $d$-dimensional polytope with $d+1$ vertices. Let $S\subset{\mathbf{N}}_{\mathbb{R}}$ be a simplex with $\operatorname{vert}(S)=\{v_{1},\cdots,v_{d+1}\}$. Given a point $x\in{\mathbf{N}}_{\mathbb{R}}$, the barycentric coordinates of $x$ with respect to $S$ are the uniquely determined numbers $\beta_{1},\cdots,\beta_{d+1}$ such that
@@
Given a $d$-dimensional polytope $P\subset{\mathbf{N}}_{\mathbb{R}}$, the normalized volume of $P$ is $\operatorname{Vol}(P)=d!\operatorname{vol}_{\mathbb{Z}}(P)$, where $\operatorname{vol}_{\mathbb{Z}}(P)$ is the (Euclidean) volume of $P$. For convenience, we call $\operatorname{Vol}(P^{*})$ (resp. $\operatorname{vol}_{\mathbb{Z}}(P^{*})$) the dual normalized volume (resp. dual volume) of $P$.
In each dimension $d$, [15, Lemma 5] gives an upper bound for the volumes of $d$-dimensional simplices containing interior lattice points in terms of the number of interior lattice points and the barycentric coordinates of a interior lattice point.
@@
Let $S$ be a $d$-dimensional simplex in ${\mathbf{N}}_{\mathbb{R}}$. Let $n=|\operatorname{int}(S)\cap{\mathbf{N}}|$ be the number of interior lattice points in $S$. Suppose $n\geq 1$ and $x\in\operatorname{int}(S)\cap{\mathbf{N}}$ has barycentric coordinates $\beta_{1}\geq\cdots\geq\beta_{d+1}$. Then
@@
Note that for a Fano simplex $S$, the dual $S^{*}$ contains exactly one lattice point ${\mathbf{0}}$ in its interior. So we have the following corollary to bound the dual volume of a Fano simplex.
@@
Let $S$ be a $d$-dimensional Fano simplex in ${\mathbf{N}}_{\mathbb{R}}$. Suppose ${\mathbf{0}}\in\operatorname{int}(S)$ has barycentric coordinates $\beta_{1}\geq\cdots\geq\beta_{d+1}$. Then
@@
To find the upper bound for anti-canonical volumes of certain toric Fano varieties with Picard number one, we need to bound the the dual normalized volumes of the corresponding Fano simplices from above (see Subsection 2.2 below). For this purpose, we need to find out some effective inequalities for barycentric coordinates of the origin ${\mathbf{0}}$. We will use the following Theorem to derive some important inequalities for barycentric coordinates of the origin ${\mathbf{0}}$ with respect to certain Fano simplices (see Theorem 3.1 below).
@@
Let $A$ be a real $d\times d$ matrix such that $10<|\det(A)|<$ 1. Then there exists an integer point $z\in{\mathbb{Z}}^{d}\setminus\{{\mathbf{0}}\}$ such that $||A\cdot z||_{\infty}<1$, where $||\cdot||_{\infty}$ denotes the $l_{\infty}$-norm associated to the vector spaces $\mathbb{R}^{d}$.
We will use the strategy in [5] to bound the dual normalized volume of an $\epsilon$-lc Fano polytope for some $\epsilon>0$.
@@
Fix a positive real number $\epsilon>0$. We say a $d$-dimensional $\epsilon$-lc Fano polytope $P\subset{\mathbf{N}}_{\mathbb{R}}$ is minimal if for each $v\in\operatorname{vert}(P)$, the convex hull $\operatorname{conv}(P\cap{\mathbf{N}}\setminus\{v\})$ is not a $d$-dimensional $\epsilon$-lc Fano polytope in ${\mathbf{N}}_{\mathbb{R}}$.
Given a $d$-dimensional $\epsilon$-lc Fano polytope $P\subset{\mathbf{N}}_{\mathbb{R}}$, we can find a $d$-dimensional minimal $\epsilon$-lc Fano polytope $P^{\prime}\subset P$ by possibly removing several vertices of $P$. Note that $P^{\prime}\subset P$ implies $\operatorname{vol}(P^{\prime*})\geq\operatorname{vol}(P^{*})$. So to find an upper bound for the dual volume of any $\epsilon$-lc Fano polytope, it is enough to consider only minimal $\epsilon$-lc Fano polytopes, in which case we will use the decomposition of a minimal $\epsilon$-lc Fano polytope into lower dimensional minimal $\epsilon$-lc Fano simplices:
Theorem 2.7 (cf. [13, Proposition 3.2] and [5, Corollary 2.3]).
Let $P\subset{\mathbf{N}}_{\mathbb{R}}$ be a $d$-dimensional minimal $\epsilon$-lc Fano polytope. If $P$ is not a simplex, then there exist lower dimensional minimal $\epsilon$-lc Fano simplices $S_{1},\cdots,S_{t}(2\leq t\leq d)$ such that $P=\operatorname{conv}(S_{1}\cup\cdots\cup S_{t})$, where $\operatorname{vert}(S_{i})\subset\operatorname{vert}(P)(1\leq i\leq t)$. For $1\leq i\leq t$, denote $d_{i}:=\dim S_{i}$ and $r_{i}:=|\operatorname{vert}(S_{i})\cap\operatorname{vert}(P_{i-1})|$, where $P_{i-1}=\operatorname{conv}(S_{1}\cup\cdots\cup S_{i-1})$ for $2\leq i\leq t$ and $P_{0}$ is the empty set. We have the following:
@@
In the proof of [13, Proposition 3.2], Only the property that Fano polytopes have ${\mathbf{0}}$ in the interior really makes a difference. So the decomposition strategy also works for minimal $\epsilon$-lc Fano polytopes.
2.2. Anti-canonical volumes of toric Fano varities
@@
Let $X=T_{{\mathbf{N}}}\operatorname{emb}(\Delta)$ be a $d$-dimensional $\epsilon$-lc toric Fano variety. Denote by $P_{X}$ the convex hull $\operatorname{conv}(e_{i},e_{i}\in\Delta(1))$, where $\Delta(1)$ is the set of primitive generators of one-dimensional rays of $\Delta$. Then
$\operatorname{mld}(X)=\inf\{t>0|tP_{X}\cap{\mathbf{N}}\supsetneq\{{\mathbf{0}}\}\}\geq\epsilon,$
@@
$P_{X}^{*}=\{m\in{\mathbf{M}}_{\mathbb{R}}|<m,e_{i}>+1\geq 0,\forall e_{i}\in\Delta(1)\}.$
According to [8, Theorem 13.4.3], we have
@@
There are known results for the optimal upper bounds of anti-canonical volumes of canonical toric Fano varieties in dimension $d\geq 3$.
Theorem 2.9 (cf. [13, Theorem 4.6] and [5, Corollary 1.3]).
@@
$\operatorname{Vol}(-K_{X})\leq 2u_{d,1}^{2},$
where equality holds if $X\simeq\mathbb{P}(\frac{2u_{d,1}^{2}}{1+u_{1,1}},\cdots,\frac{2u_{d,1}^{2}}{1+u_{d-1,1}},1,1)$.

### v2

@@ 1.1. Upper bounds for anti-canonical volumes of Fano varieties
and for $q\geq 2$, the equality holds if and only if
$X\simeq\mathbb{P}(\frac{2u_{d,q}}{1+u_{1,q}},\cdots,\frac{2u_{d,q}}{1+u_{d-1,q}},1,1).$
@@
1.2. Upper bounds for dual normalized volumes of Fano polytopes
@@
Fix two integers $d\geq 3$ and $q\geq 2$. Let $P$ be a $d$-dimensional $\frac{1}{q}$-lc Fano polytope. Then
@@
and the equality holds if and only if
$P^{*}\simeq\operatorname{conv}(\frac{1+u_{1,q}}{q}e_{1},\cdots,\frac{1+u_{d-1,q}}{q}e_{d-1},\mp\frac{u_{d,q}}{q}e_{d}),$
where $e_{i}$ are the standard basis of $\mathbb{R}^{d}$.
@@ 1.3. Upper bounds for volumes of lattice simplices
At the same time, we can get the following optimal upper bound for the volumes of lattice simplices $S\subset\mathbb{R}^{d}$ such that $\operatorname{int}(\frac{1}{q}S)\cap{\mathbb{Z}}^{d}=\{{\mathbf{0}}\}$.
@@
Fix two integers $d\geq 3$ and $q>0$. Let $S$ be a $d$-dimensional latttice simplex in $\mathbb{R}^{d}$ such that $\operatorname{int}(\frac{1}{q}S)\cap{\mathbb{Z}}^{d}=\{{\mathbf{0}}\}$. Then
$\operatorname{vol}_{{\mathbb{Z}}}(S)\leq\frac{2u_{d,q}^{2}}{d!q},$
@@ 1.4. Organization of this paper
In Section 2, we collect the necessary material from geometry of numbers and toric geometry, introducing the relevant polytopes, the way to bound the volume of a simplex (see Theorem 2.3) and the decomposition strategy to deal with the dual volume of certain polytopes (see Theorem 2.7); In Section 3, we introduce generalized Product-Sum inequalities (see Thereom 3.1) for barycentric coordinates of the origin ${\mathbf{0}}$ with respect to an $\frac{1}{q}$-lc Fano simplex and use it to characterize when the products of barycentric coordinates attain the minimal ones (see Theorems 3.8 and 3.9); In Section 4, based on the results in Section 3, we apply Corollary 2.4 to get an upper bound for the dual normalized volume of any $d$-dimensional $\frac{1}{q}$-lc Fano simplex when $d\geq 2$, and hence an upper bound for the anti-canonical volumes of $d$-dimensional $\frac{1}{q}$-lc toric Fano varieties of Picard number one (see Theorem 4.1 and Theorem 4.2). We construct examples to show that these upper bounds are optimal when $d\geq 3$ (see Example 4.3), and we also show that under some conditions these examples are the unique ones which attain the optimal upper bounds; In Section 5, we use the decomposition strategy and integration method from [6] to find upper bounds for the dual normalized volumes of $\frac{1}{q}$-lc Fano polytopes which are not simplices; In Section 6, we prove Theorem 1.1, Theorem 1.2 and Theorem 1.3.
@@ 2.1. Geometry of numbers
We refer to [11] for the basic knowledge in geometry of numbers. Fix a positive integer $d$. Let $\mathbb{R}^{d}$ be the $d$-dimensional vector space and ${\mathbb{Z}}^{d}\subset\mathbb{R}^{d}$ the lattice of rank $d$. A unimodular transformation is an affine bijection map $\phi:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{d}$ such that $\phi({\mathbb{Z}}^{d})={\mathbb{Z}}^{d}$. A polytope $P$ is a convex bounded subset of $\mathbb{R}^{d}$ such that $P$ is the intersection of finitely many affine half-spaces. Two polytopes $P,Q$ are said to be unimodularly equivalent, denoted by $P\simeq Q$, if there exists a unimodular transformation $\phi$ such that $\phi(P)=Q$. For a polytope $P$, we use $\operatorname{int}(P)$ and $\operatorname{vert}(P)$ to denote the interior of $P$ and the set of vertices of $P$ respectively. The dual of a polytope $P$ is defined as
$P^{*}:=\{m\in\mathbb{R}^{d}|<m,v>+1\geq 0,\forall v\in P\}.$
@@
A polytope $P\subset\mathbb{R}^{d}$ is called a lattice polytope if $\operatorname{vert}(P)\subset{\mathbb{Z}}^{d}$. We say $P\subset\mathbb{R}^{d}$ is a Fano polytope if it is a lattice polytope with ${\mathbf{0}}\in\operatorname{int}(P)$ and each vertex $v\in\operatorname{vert}(P)$ is a primitive lattice point. A Fano polytope $P$ is called an $\epsilon$-lc Fano polytope for some $\epsilon>0$ if $\operatorname{int}(\epsilon P)\cap{\mathbb{Z}}^{d}=\{{\mathbf{0}}\}$. Here ${\mathbf{0}}$ denotes the origin of $\mathbb{R}^{d}$.
@@
A simplex $S\subset\mathbb{R}^{d}$ is a $d$-dimensional polytope with $d+1$ vertices. Let $S\subset\mathbb{R}^{d}$ be a simplex with $\operatorname{vert}(S)=\{v_{1},\cdots,v_{d+1}\}$. Given a point $x\in\mathbb{R}^{d}$, the barycentric coordinates of $x$ with respect to $S$ are the uniquely determined numbers $\beta_{1},\cdots,\beta_{d+1}$ such that
@@
Given a $d$-dimensional polytope $P\subset\mathbb{R}^{d}$, the normalized volume of $P$ is $\operatorname{Vol}(P)=d!\operatorname{vol}_{\mathbb{Z}}(P)$, where $\operatorname{vol}_{\mathbb{Z}}(P)$ is the (Euclidean) volume of $P$. For convenience, we call $\operatorname{Vol}(P^{*})$ (resp. $\operatorname{vol}_{\mathbb{Z}}(P^{*})$) the dual normalized volume (resp. dual volume) of $P$.
In each dimension $d$, [16, Lemma 5] gives an upper bound for the volumes of $d$-dimensional simplices containing interior lattice points in terms of the number of interior lattice points and the barycentric coordinates of an interior lattice point.
@@
Let $S$ be a $d$-dimensional simplex in $\mathbb{R}^{d}$. Let $n=|\operatorname{int}(S)\cap{\mathbb{Z}}^{d}|$ be the number of interior lattice points in $S$. Suppose $n\geq 1$ and $x\in\operatorname{int}(S)\cap{\mathbb{Z}}^{d}$ has barycentric coordinates $\beta_{1}\geq\cdots\geq\beta_{d+1}$. Then
@@
Note that for any $d$-dimensional lattice polytope $P\subset\mathbb{R}^{d}$ with ${\mathbf{0}}\in\operatorname{int}(P)$, we have $\operatorname{int}(P^{*})\cap{\mathbb{Z}}^{d}=\{{\mathbf{0}}\}$. In particular, for a Fano simplex $S$, the dual $S^{*}$ contains exactly one lattice point ${\mathbf{0}}$ in its interior. So we have the following corollary to bound the dual volume of a Fano simplex.
@@
Let $S$ be a $d$-dimensional Fano simplex in $\mathbb{R}^{d}$. Suppose ${\mathbf{0}}\in\operatorname{int}(S)$ has barycentric coordinates $\beta_{1}\geq\cdots\geq\beta_{d+1}$. Then
@@
Proof.
This follows from Theorem 2.3 and the fact that the barycentric coordinates of ${\mathbf{0}}$ with respect to $S^{*}$ and $S$ are the same. ∎
To find the upper bound for anti-canonical volumes of certain toric Fano varieties, we need to bound the dual normalized volumes of the corresponding Fano polytopes from above (see Subsection 2.2 below). Our first step is to bound the dual normalized volumes of certain Fano simplices. As indicated by Corollary 2.4, it suffices to bound the product of the first $d$ barycentric coordinates of the origin ${\mathbf{0}}$ from below. For this purpose, we need to find out some effective inequalities for these barycentric coordinates.
We will use the following Theorem to derive some important inequalities for barycentric coordinates of the origin ${\mathbf{0}}$ with respect to certain Fano simplices (see Theorem 3.1 below).
@@
Let $A$ be a real $d\times d$ matrix such that $10<|\det(A)|<$ 1. Then there exists an integer point $z\in{\mathbb{Z}}^{d}\setminus\{{\mathbf{0}}\}$ such that $||A\cdot z||_{\infty}<1$, where $||\cdot||_{\infty}$ denotes the $l_{\infty}$-norm associated to the vector space $\mathbb{R}^{d}$.
To bound the dual normalized volumes of Fano polytopes which are not simplices, we will use the strategy in [6]. Therefore, we need to introduce the concept of minimal $\epsilon$-lc Fano polytopes.
@@
Fix a positive real number $\epsilon>0$. We say a $d$-dimensional $\epsilon$-lc Fano polytope $P\subset\mathbb{R}^{d}$ is minimal if for each $v\in\operatorname{vert}(P)$, the convex hull $\operatorname{conv}(P\cap{\mathbf{N}}\setminus\{v\})$ is not a $d$-dimensional $\epsilon$-lc Fano polytope in $\mathbb{R}^{d}$.
Note that by definition, every $\epsilon$-lc Fano simplex is minimal.
Given a $d$-dimensional $\epsilon$-lc Fano polytope $P\subset\mathbb{R}^{d}$, we can find a $d$-dimensional minimal $\epsilon$-lc Fano polytope $P^{\prime}\subset P$ by possibly removing several vertices of $P$. Note that $P^{\prime}\subset P$ implies $\operatorname{vol}_{{\mathbb{Z}}}(P^{\prime*})\geq\operatorname{vol}_{{\mathbb{Z}}}(P^{*})$. So to find an upper bound for the dual volume of any $\epsilon$-lc Fano polytope, it is enough to bound the dual volumes for minimal $\epsilon$-lc Fano polytopes from above, in which case we will use the decomposition of a minimal $\epsilon$-lc Fano polytope into lower dimensional $\epsilon$-lc Fano simplices:
Theorem 2.7 (cf. [14, Proposition 3.2] and [6, Corollary 2.3]).
Let $P\subset\mathbb{R}^{d}$ be a $d$-dimensional minimal $\epsilon$-lc Fano polytope. If $P$ is not a simplex, then there exist lower dimensional $\epsilon$-lc Fano simplices $S_{1},\cdots,S_{t}(2\leq t\leq d)$ such that $P=\operatorname{conv}(S_{1}\cup\cdots\cup S_{t})$, where $\operatorname{vert}(S_{i})\subset\operatorname{vert}(P)(1\leq i\leq t)$. For $1\leq i\leq t$, denote $d_{i}:=\dim S_{i}$ and $r_{i}:=|\operatorname{vert}(S_{i})\cap\operatorname{vert}(P_{i-1})|$, where $P_{i-1}=\operatorname{conv}(S_{1}\cup\cdots\cup S_{i-1})$ for $2\leq i\leq t$ and $P_{0}$ is the empty set. We have the following:
@@
In the proof of [14, Proposition 3.2], only the property that Fano polytopes have ${\mathbf{0}}$ in the interior really makes a difference. So the decomposition strategy also works for minimal $\epsilon$-lc Fano polytopes.
2.2. Anti-canonical volumes of toric Fano varieties
@@
Let $X$ be a $d$-dimensional $\epsilon$-lc toric Fano variety associated to the fan $\Delta\subset\mathbb{R}^{d}$. Denote by $P_{X}$ the convex hull $\operatorname{conv}(e_{i},e_{i}\in\Delta(1))$, where $\Delta(1)$ is the set of primitive generators of one-dimensional rays of $\Delta$. Then
$\operatorname{mld}(X)=\inf\{t>0|tP_{X}\cap{\mathbb{Z}}^{d}\supsetneq\{{\mathbf{0}}\}\}\geq\epsilon,$
@@
$P_{X}^{*}=\{m\in\mathbb{R}^{d}|<m,e_{i}>+1\geq 0,\forall e_{i}\in\Delta(1)\}.$
According to [9, Theorem 13.4.3], we have
@@
The following are known results for the optimal upper bounds for anti-canonical volumes of canonical toric Fano varieties in dimension $d\geq 3$.
Theorem 2.9 (cf. [14, Theorem 4.6], [3, Theorem 2.11] and [6, Corollary 1.3]).
@@
$\operatorname{Vol}(-K_{X})\leq 2u_{d,1}^{2}.$
For $d=3$, the equality holds if and only if $X\simeq\mathbb{P}(6,4,1,1)$ or $\mathbb{P}(3,1,1,1)$; for $d\geq 4$, the equality holds if and only if $X\simeq\mathbb{P}(\frac{2u_{d,1}}{1+u_{1,1}},\cdots,\frac{2u_{d,1}}{1+u_{d-1,1}},1,1)$.